# Amorphous Boron Nitride as an Ultrathin Copper Diffusion Barrier for Advanced Interconnects

### v3

@@ 3.2 Mechanical Properties
Having good mechanical properties is crucial for diffusion barriers, as low mechanical strength and Young’s modulus complicate the integration process. Moreover, good mechanical properties indicate that the film can resist deformation under stress which is crucial for structural integrity during the operation [2, 38, 39]. As other properties, mechanical properties are strongly influenced by the morphological features and fabrication process of -BN films. Here, we investigated the Young’s, bulk and shear moduli of -BN films by computing the elastic constants as a function of film thickness and cooling rate.
@@
The elastic constants of each film were calculated as explained in the Methods section. Young’s modulus and other mechanical moduli of films respect to the thickness and applied cooling rate is presented in Fig. 4. The resulting mechanical properties exhibit trends that parallel the morphological features. Films generated with higher cooling rates, which as seen above present reduced numbers of - and -hybridized atoms, an increased number of homonuclear bonds and higher porosity, display a lower Young’s modulus compared to samples that were cooled slower. Other moduli values are also decreased significantly with higher cooling rates. High Young’s modulus of slow-cooled samples can attributed to high number of - and -hybridized atoms and absence of pores and voids that we observed in fast-cooled samples. The positive impact of -hybridized atoms on mechanical properties is well-established in the literature [18, 19, 40]. Young’s modulus is particularly critical for diffusion barriers, as a higher value indicates greater stiffness and resistance to deformation. This is essential for ensuring that the barrier maintains its structural integrity during the high-stress conditions typical in semiconductor integration and operation. A higher Young’s modulus in slow-cooled films suggests that these barriers would be better at resisting compressive and tensile stresses, minimizing risks such as cracking or delamination that could compromise the diffusion barrier’s effectiveness. In contrast, the lower Young’s modulus observed in fast-cooled films may render them more susceptible to deformation under stress, which could lead to voids and defects during device fabrication and operation. However, it should be noted that high Young’s modulus in barrier layers could be problematic if there is a mismatch between the thermal expansion coefficients of the barrier and other layers. Since high Young’s modulus leads to more stiff structure, such a mismatch could lead to increased stress and structural deformations.
@@
Although not explicitly shown in the figure, Poisson’s ratio exhibits a weak dependence on applied cooling rates. For films cooled at 5 K/ps and 10 K/ps, Poisson’s ratio remains nearly constant at approximately 0.28–0.30 across different thicknesses, suggesting a consistent balance between lateral expansion and contraction under stress. In contrast, faster-cooled samples exhibit a noticeably lower Poisson’s ratio (0.19–0.22), suggesting a stiffer response to deformation. This change, though subtle, may affect the barrier’s ability to adapt to mechanical strains and its performance during thermal cycling and integration processes.
@@ 3.3 Dielectric Properties of -BN Samples
The use of a low dielectric constant barrier would be desirable compared to current metal barriers, as the absence of a metal barrier increases the cross-sectional area available for low-resistivity metals, such as Cu or other metals with better electromigration characteristics. -BN has already been studied for this use [28, 15, 12], however both experimental and theoretical results show a wide range dielectric constant values, strongly suggesting that this property, as well as mechanical and structural properties, depends on the morphology.
@@
Fig. 5 shows the z-component of the real and imaginary parts of the dielectric function for 3-nm (a, c) and 5-nm thick (b, d) films.555Electronic properties for 7-nm films were not evaluated due to computational limitations. The results are quite similar at both thicknesses. While slow-cooled samples (CR25 K/ps) at both thicknesses exhibited relatively low dielectric constants, around 2.8 (Fig. 5-a), the dielectric functions of the two fastest-cooled samples displayed a tendency to diverge at low frequencies due to the absence of an electronic gap. Although there seems to be a relationship between cooling rates and dielectric constants, the values are too close to establish a definitive trend given the precision of the model.
@@
To better understand these effects, it is of interest to consider the energy and oscillator strengths of electronic transitions in the sample. Indeed, , up to constants, can be seen as the sum of oscillator strengths over all transition the system, inversely weighted by the transition energies (see equation 1 and, e.g. [28]). Conveniently, the imaginary part of the dielectric function resolves in energy the density and oscillator strengths of electronic transitions (as can be seen from the delta function in equation 2). For fast-cooled samples, it displays peaks at low energies, signaling the existence of low-energy transitions (of non-vanishing oscillator strengths), which are absent in the slowly-cooled samples.
@@
To get further insight into these transitions and the dielectric constant behaviours, we examine the DOS and IPR of the films. In Fig. 6, we present, for simplicity, the 3-nm thick films generated with cooling rates of 5 K/ps and 100 K/ps. Both films can be seen to exhibit mid-gap states. However, while the slow-cooled sample (6-a) shows a nominal (including mig-gap states) electronic gap of eV, the fast-cooled sample (6-b) displays an effectively closed gap, allowing for the aforementioned low-energy transitions. In the slow-cooled sample, it appears that most midgap states are unoccupied (red dots in Fig. (6-a)), while the fast-cooled sample exhibits occupied mid-gap states. In both cases, mid-gap states show IPRs0, indicating real-space localization, which typically suppresses momentum matrix elements between them (numerators in equations 1 and 2) and DC conductivity, lowering (but here not eliminating) their impact on the dielectric constant.
@@
At this point, we must recall the limitations of our electronic model and give further qualitative remarks. In relation to midgap states, DFT calculations in [28] suggest that the dielectric constant of bulk -BN is increased by the presence of - and - coordinated atoms (in -dominated samples) and low short range order, which can be related to the presence of homonuclear (B-B and N-N) bonds. Lin et al. [15] also point towards the detrimental effects of Boron clusters. Similarly, Glavin et al. [12] indicated that B-B bonds may promote the formation of mid-gap states or conductive pathways in thin films. These remarks should be put in regard with the findings of section 3.1. We further note that although the sample cooled at 25 K/ps had a significant number of N-N bonds, we did not observe any large N or B clusters. However, we observed some B and N clusters (B-B or N-N bonds with a size of 4-12 atoms). The aforementioned morphological features, which we expect in fast-cooled films; especially in the case of thinner films (see Fig. 3) may contribute to the formation of midgap states and significantly affect the dielectric constant in these films. To a degree, we observe such a trend in the electronic calculations presented here, although the simplicity of the tight-binding model used here may underestimate these effects, making quantitative statements difficult. Finally, we should comment on the effects of density: quickly-cooled films; especially the thinnest ones, present a lower density (Fig. 3 (d)), which favours a lower dielectric constant (see the presence of the unit cell volume in the denominator of equation 1). However, this is accompanied by an increase in undesirable morphological features ( fraction, homonuclear bonds). Within the caveats above, our results suggest that the net effect is not favourable for ultralow dielectric constant properties.
@@ 4 Theoretical Investigation of -BN Diffusion Barriers
Fig. 7 presents the Cu diffusivity as a function of the inverse of temperature within 3-nm (a), 5-nm (b) and 7-nm (c) thick films. We assume that diffusivity of Cu atoms should be negligible as long as the -BN barrier can prevent the Cu diffusion.
Note that, due to the effects of temperature on atomic motion, diffusivity is expected to increase with temperature (as seen in Fig.7-c), so that only a sharp increase of diffusivity (as seen in Fig.7-b and -c) can be safely interpreted as a sign of barrier failure. All samples, regardless the thickness and morphology, were able to prevent the Cu diffusion at low temperatures. 7-nm thick films prevent or largely mitigate Cu diffusion even at high temperatures (up to 1000 K). In contrast, 3-nm and 5-nm thick films were only successful when lower cooling rates were applied during their formation.
As with all properties, the barrier performance of films largely depends on their morphology. Thus, we focused on the morphology of films to better understand the trends shown in Fig. 7. In the case of 7-nm thick films, even films formed with the highest high cooling rates exhibited relatively low levels of -hybridized atoms, high density, and fewer nanovoids near the boundaries compared to thinner films. While films cooled rapidly (CR25 K/ps) have some voids and -hybridized atoms leading to very limited Cu diffusion at near interface, slowly cooled films had a smooth surface without any voids close to the interface, therefore no Cu diffusion. The consistent increase in Cu diffusivity is largely due to increased atomic motion as the temperature increases. The superior diffusion barrier performance of 7-nm films can be attributed to their reduced defects and increased structural order, as described in Section 3.1, which allows them to better resist atomic migration even under less favourable conditions.
In contrast, the morphology of 3-nm and 5-nm thick films is significantly different than the 7-nm thick films. At higher cooling rates, the proportion of -hybridized bonds increased rapidly and voids were observed near the interface between -BN and Cu films. Moreover, the surface of rapidly cooled 3-nm and 5-nm thick samples were noticeably rough. The potential energy per atom shows that these films are less stable than the 7-nm thick samples at the same cooling rates, as shown in Fig.12 in B, which can be attributed to larger number of -hybridized atoms and homonuclear bonds in thinner films. The lower density at high cooling rates (1.9-2.0 for 3-nm and 5-nm thick films, for 7-nm thick films at CR25 K/ps) along with the other morphological features mentioned here, indicates the existence of the voids in the films cooled rapidly. These voids and -hybridized atoms created pathways through the barrier which allowed Cu atoms to diffuse into the -BN. While 3-nm and 5-nm films are more susceptible to these defects, films generated at lower cooling rates demonstrated morphological stability, showing that quality, rather than thickness alone, determines performance. As Cu atoms close to the interface diffused into the -BN layer, B and N atoms also migrated into the Cu layer. This process accelerated the void formation and finally -BN barriers lost their structural integrity completely. Although providing visual representation of existing voids close to the interface is not easy, we present two 3-nm films which cooled at a rate of 100 K/ps (a) and 5 K/ps (b) in Fig. 8 after the annealing process. The black line shows the initial position of the interface. As shown in the Fig. 8, after the annealing process, the slow-cooled -BN barrier stopped Cu atoms from diffusing into the structure (only a few seemed to be diffused into the near interface area, which can be considered negligible). On the other hand, the fast-cooled -BN structure contained many small voids and had a worse stability. After the annealing at high temperatures, many Cu atoms diffused into the barrier, causing it to lose its structural integrity completely. Additionally, some B and N atoms also diffused into the Cu film, although in much smaller quantity compared to the diffused Cu atoms into the -BN film. While B and N atoms largely diffused into the regions close to the interface, Cu atoms diffused even deep into the -BN. We observed a similar process in other fast-cooled samples (CR: 25-100 K/ps) and 5-nm thick samples as well. However, it should be noted that 5-nm thick films managed to stop Cu diffusion at higher temperatures, however, fast-cooled 5-nm thick films were not able to prevent the Cu diffusion.
These differences in morphology, caused by the different cooling protocols, demonstrate that the barrier performance of the -BN films largely depends on the film quality. While 3- and 5-nm thick films generated with cooling rates of 5 K/ps and 10 K/ps could stop the Cu diffusion until 1000 K, films with high coordination defects and voids (cooling rates greater than 50 K/ps) only managed to do the same until 500-600 K. These results suggest that ultrathin -BN films, when generated under optimised conditions, can perform comparably to thicker films as effective diffusion barriers and dielectric layers in integrated circuits using Cu wiring.

### v4

@@ 3.2 Mechanical Properties
Strong mechanical performance is critical for diffusion barriers because low strength and a low Young’s modulus complicate integration. A robust mechanical response indicates that the film resists deformation under stress, which is crucial for structural integrity during processing and operation [2, 41, 42]. As with other properties, the mechanical behaviour of -BN films is strongly influenced by morphology and fabrication conditions. Here, we investigated the Young’s, bulk, and shear moduli of -BN films by computing the elastic constants as a function of film thickness and cooling rate.
@@
The elastic constants for each film were calculated as described in the Methods section. Young’s modulus and the other mechanical moduli with respect to thickness and applied cooling rate are presented in Fig. 4. The resulting mechanical properties exhibit trends that parallel the morphological features. Films generated at higher cooling rates, as presented in Fig. 3, contain reduced fractions of - and -hybridised atoms, a higher proportion of homonuclear bonds, lower density, and increased porosity (Fig. 13 in B). These films display a lower Young’s modulus than more slowly cooled samples. The other moduli also decrease significantly with increasing cooling rate. The high Young’s modulus of the slow-cooled samples can be attributed to the higher fractions of - and -hybridised atoms and to the absence of the pores and voids observed in the fast-cooled samples (Fig. 13). The positive impact of -hybridised environments on stiffness is well established [18, 19, 43]. Young’s modulus is particularly critical for diffusion barriers because a higher value indicates greater stiffness and resistance to deformation. This is essential to ensure that the barrier maintains structural integrity during the high-stress conditions typical of semiconductor processing and operation. A higher Young’s modulus in the slow-cooled films suggests that these barriers would better resist compressive and tensile stresses, minimising risks such as cracking or delamination that could compromise diffusion-barrier performance. In contrast, the lower Young’s modulus observed in the fast-cooled films may render them more susceptible to deformation under stress, which could lead to voids and defects during device fabrication and operation. However, a very high Young’s modulus in the barrier layer can be problematic if there is a mismatch between the thermal expansion of the barrier and adjacent layers, since a stiffer structure can concentrate stress and promote deformation.
@@
Although not shown in the figure, Poisson’s ratio exhibits a weak dependence on the applied cooling rate. For films cooled at 5 K/ps and 10 K/ps, Poisson’s ratio remains approximately constant at 0.28–0.30 across thicknesses, indicating a consistent balance between lateral expansion and contraction under stress. In contrast, faster-cooled samples exhibit a lower Poisson’s ratio (0.19–0.22), indicating reduced lateral compliance. This subtle change may affect the barrier’s ability to accommodate mechanical strains and its performance during thermal cycling and integration processes.
Additionally, the mechanical moduli exhibit a clear dependence on film thickness, thinner films consistently exhibit lower values across all cooling rates. This behaviour is consistent with the morphological differences discussed in 3.1. Specifically, thinner films contain a higher fraction of -hybridised atoms, a larger proportion of homonuclear bonding, lower mass density, and increased porosity (see Fig. 3 and B). These structural features reflect a higher degree of disorder and reduced network connectivity, resulting in a softer and mechanically less robust amorphous matrix. In contrast, enhanced structural relaxation in thicker films allows them to accommodate local strain and form a more rigid network, which increases stiffness.
@@ 3.3 Dielectric Properties of -BN Samples
Replacing the current metal barriers with low dielectric constant material is desirable because it increases the the cross-sectional area available for low-resistivity metals (such as Cu) and reduces the overall line resistivity. -BN has already been studied for this use [31, 15, 12], however both experimental and theoretical results show a wide range dielectric constant values, strongly suggesting that this property, as well as mechanical and structural properties, depends on the morphology [44].
@@
Fig. 5 shows the z-component of the real and imaginary parts of the dielectric function for 3-nm (a, c) and 5-nm thick (b, d) films.444Electronic properties for 7-nm films were not evaluated due to computational limitations. The results are quite similar at both thicknesses. While slow-cooled samples (CR25 K/ps) at both thicknesses exhibited relatively low dielectric constants, around 2.8 (Fig. 5-a), the dielectric functions of the two fastest-cooled samples displayed a tendency to diverge at low frequencies due to the absence of an electronic gap. Although there seems to be a relationship between cooling rates and dielectric constants, the values are too close to establish a definitive trend of the variability of the dielectric properties upon increasing the film thickness.
@@
To further analyse the obtained results, it is of interest to consider the energy and “strengths” of electronic transitions in the sample.555Up to constant prefactors, the are essentially oscillator strengths along divided by transition energies. Indeed, , up to constants, can be seen as the sum of these strengths over all transitions the system, inversely weighted by the transition energies (see equation 1 and, e.g. [31]). Conveniently, the imaginary part of the dielectric function resolves in energy the density and strengths of electronic transitions (per equation 2, it can be interpreted as a density of states for electronic transitions, with each transition weighted by the corresponding strength [32]). For fast-cooled samples, the imaginary part of the dielectric function displays peaks at low energies, signaling the existence of low-energy transitions (of non-vanishing strengths), which are absent in the slowly-cooled samples.
@@
To get further insight into these transitions and the dielectric constant behaviour, we examine the DOS and IPR of the films. In Fig. 6, we present, for simplicity, the 3-nm thick films generated with cooling rates of 5 K/ps and 100 K/ps. Both films can be seen to exhibit mid-gap states. However, while the slow-cooled sample (6-a) shows a nominal (including mig-gap states) electronic gap of eV, the fast-cooled sample (6-b) displays an effectively closed gap, allowing for the aforementioned low-energy transitions. In the slow-cooled sample, it appears that most midgap states are unoccupied (red dots in Fig. (6-a)), while the fast-cooled sample exhibits occupied mid-gap states. In both cases, mid-gap states show IPRs0, indicating real-space localisation, which typically suppresses momentum matrix elements between them (numerators in equations 1 and 2) and DC conductivity, lowering (but here not eliminating) their impact on the dielectric constant.
@@
At this point, we should provide qualitative remarks regarding the expected impacts of structure on the dielectric constant of aBN systems. DFT calculations in [31] suggest that the dielectric constant of bulk -BN is increased by the presence of - and - coordinated atoms (in -dominated samples) and local structural disorder, which can be related to the presence of homonuclear (B-B and N-N) bonds. Lin et al. [15] also point towards the detrimental effects of Boron clusters.666Although the 3 nm thick sample cooled at 25 K/ps had a significant number of N-N bonds, we did not observe any large N or B clusters. However, we observed some B and N clusters (B-B or N-N bonds with a size of 4-12 atoms). Similarly, Glavin et al. [12] indicated that B-B bonds may promote the formation of mid-gap states or conductive pathways in thin films. These features, which we expect in fast-cooled films, as shown in Fig. 3, may thus contribute to the formation of midgap states and increase the dielectric constant in these films.
As can be inferred from the presence of the unit cell volume in the denominator of equation 1, with all other parameters unchanged, the dielectric constant is also expected to increase with density. Our results (Fig. 3 d.) show that density decreases in fast-cooled films compared to slower cooling rates. This suggests an opposite trend to the one discussed above: from this effect alone, we would expect a decrease of on the order of from the 3-nm and 5-nm films cooled at compared to those cooled at (as per the changes in density reported in Fig. 3 d.). As can be seen in Fig. 5, we do not observe such a reduction, which hints at the importance of the aforementioned structure-induced changes to the electronic structure.
Thus, while fast cooling rates favor a lower density (and thus volume), which, if unit cell polarizability was unchanged, would lower the dielectric constant, it also favors morphological features (homonuclear bonds, - and - coordinated atoms, etc.) which can increase the polarizability per unit cell.
It is possible that, within our simple tight-binding model, these effects largely cancel each other, leading only to a very weak dependence of dielectric constant on cooling rates. As stated above, we still observe a very weak effect for both thicknesses suggesting that could increase with faster cooling rates. Considering the simplicity of the tight-binding model, we cannot however confidently establish a definite trend. Within this caveat, we nevertheless note that thinner (3 nm) films appear to display more significant variations of their dielectric functions (and constants) with cooling rate. This appears consistent with the findings of section 3.1: the thinner the film, the stronger the effects of cooling rate variations in inducing homonuclear bonds and - and - coordinated atoms.
From Fig. 6 we observe that there appears to be a gap of between unoccupied and occupied states with low IPR (thus expected to be more delocalised with respect to the midgap states). Observation of , in Fig. 5, suggests that this is the case in most samples, with these states giving rise to the first main peak (at the single-particle level), albeit lower energy contributions from the (significantly less numerous) mid-gap states can also be observed. Here again, the strongest change with cooling rate is observed for the 3-nm films. For this thickness, in the case of the cooling rate, we observe a significantly higher response at low () frequencies (Fig 5 b.), which correspondingly increases the dielectric constant (Fig 5 a.). This effect is significantly less pronounced in the case of 5-nm films, further hinting at the fact that properties of thicker films may be less dependent on cooling rate.
@@ 4 Theoretical Investigation of -BN Diffusion Barriers
Fig. 7 presents the Cu diffusivity, , as a function of inverse temperature for 3 nm (a), 5 nm (b), and 7 nm (c) films to identify when the films lose their stability and protective abilities. values are non-zero since the atoms move slightly due to temperature. With higher temperature, we observe a linear increase in log of Cu as expected. While behaviour up to 450 K is linear-like at all thickness and cooling rates, we start to observe some deviations from Arrhenius behaviour with increasing temperature. Up to this temperature, we assume that the -BN barriers are structurally stable and do not allow Cu atoms diffuse into the structure. However, with higher temperatures, amorphous films with higher bond and coordination defect concentration (B-B/N-N bonds, -hybridised atoms) and lower density (which indicates larger free volume and nanovoids) may suffer from structural changes [45] with higher temperatures and diffusion of Cu atoms. The deviation from linear Arrhenius behaviour of Cu diffusion indicates a change in the barrier integrity and film morphology. Fig. 7 and 15 in C show that the barrier performance is strictly about the morphology of films. Beyond the temperature-dependent trends in Fig. 7, a clear structure–property link is evident between the film mass density at 300 K and copper mobility. Within each thickness subset, log decreases monotonically as density increases, consistent with reduced free volume and fewer nanovoids limiting diffusion paths as shown in Fig. 15 in C. The effect is most pronounced in thinner films, and weakens as thickness increases, which aligns with the coordination and density trends in Fig. 3-d.
Slow-cooled (5 and 10 K/ps) 3 nm thick films have lower fractions of homonuclear bonds and -hybridised atoms and have higher mass density, as discussed in Section 3.1 and Fig. 3. In addition, Fig. 14 in B shows that these films are more relaxed and stable than those generated at higher cooling rates, consistent with their higher mass density and the correspondingly lower free volume and void population compared to fast-cooled samples (CR 25 K/ps). As shown in Fig. 7, Cu atoms adjacent to these films exhibit significantly low diffusivity and follow an Arrhenius-like trend up to high temperatures (800 K), with only minor deviations beyond this point. This suggests that the structure remains largely stable throughout the annealing process and continues to prevent or strongly limit Cu diffusion. However, fast-cooled samples (25 K/ps) show a different behaviour. While films at 100 K/ps only manage to stop Cu atoms until 450 K, the ones at 50 K/ps and 25 K/ps manage to show linear-like behaviour until 600 to 700 K as shown in Fig. 7. As discussed in Section 3.1, these films contain higher fraction of bonding and coordination defects—namely homonuclear bonds and -hybridised atoms—which lower their mass density and increase free volume. With rising temperature and Cu diffusion, existing nanovoids grow and these defect-rich regions lose stability, opening pathways for Cu transport. As Cu penetrates the -BN layer near the interface, B and N atoms also migrate into the Cu layer, further accelerating void growth and ultimately leading to complete loss of barrier integrity. The difference in structural response and barrier performance can be directly observed in Fig. 8. The film generated at 100 K/ps (Fig. 8a) fails to block Cu, allowing full penetration into the -BN layer, whereas the film at 5 K/ps (Fig. 8b) successfully prevents Cu diffusion.
Similar morphology-barrier performance relationship is observed for 5 nm thick films with slightly more resilience. Films generated at 5 and 10 K/ps contain low fraction of and homonuclear bonds, have higher density, and less free volumes. These morphological features lead to more stable barrier performance that manage to stop Cu diffusion until high temperatures and Cu diffisuvity only exhibit a minimal deviation from linear Arrhenius behaviour as shown in Fig. 7 and 15 in C. However fast cooled samples generated at CR above 25 K/ps contained less stable regions with -hybridised atoms and homonuclear bonds. This result in a loss of barrier integrity and quick increase in of Cu atoms at 500 K for films generated at 100 K/ps and 700 - 800 K for films generated at 25 and 50 K/ps. Similar to 3 nm films, as temperature increases further, defect-rich regions soften and voids grow, allowing gradual Cu penetration and some B and N migration into the Cu layer. Thus, while the same mechanism seen in the 3 nm case operates here, the increased thickness raises the temperature window over which the barrier can maintain its protective function.
At 7 nm, -BN barriers exhibit a different response from the thinner films. As shown in Fig. 7-c, Cu atoms have remarkably lower diffusivity values in 7 nm systems across full temperature range. While Cu atoms show linear Arrhenius-like behaviour until 600 K, films generated at higher CR started to slightly deviate from the linear behaviour, indicating a small Cu diffusion into -BN films and small possible change in the morphology of the films. This different behaviour can be explained the morphology of the films presented in Fig. 3. Figures 3 and 14 show that 7 nm films consistently contain fewer homonuclear bonds and a lower fraction of -hybridised atoms than the thinner samples, together with higher coordination numbers, higher mass density, and correspondingly reduced free volume and voids. Even the rapidly quenched films start from a more relaxed energy state than 3 or 5 nm counterparts, which prevents early softening of weak regions during heating. These combined features demonstrate that the thicker films dilute the influence of interfacial defects and stabilise the amorphous network.
These results show that the diffusion-barrier response of -BN is governed by morphology rather than thickness alone. Slow-cooled 3–5 nm films (5–10 K/ps) maintain low Cu mobility up to 1000 K, whereas rapidly quenched networks (50 K/ps) with higher coordination defects and voids lose integrity by 500–600 K. Within each thickness, higher density together with reduced homonuclear bonding and a lower fraction of sp1-hybridised atoms lead to lower Cu diffusion. Increasing thickness mitigates, but does not remove, the sensitivity to the morphology. Practically, ultrathin -BN can match the performance of thicker layers when the growth route yields a dense, well-connected network with minimal defects and free volume.